# A -ray-emitting blazar B3 1239+376 at = 3.82 identified in a multi-wavelength context

### v1

@@ II.2 Chandra X-ray Observatory data
B3 1239+376 was observed twice by the Advanced CCD Imaging Spectrometer (ACIS, Garmire etĀ al. 2003), on 2007 Mar. 10 (ACIS-7, ObsID: 7871, PI.: C. Cheung) and 2022 Dec. 31 (ACIS-67, ObsID 26906; PI: E. Meyer). The corresponding exposure times are 5.2 and 20.1 ks, respectively. The observations were performed on the pointing mode, events were recorded in the Faint telemetry format and with the Timed Exposure mode. Data reductions were carried out using the Chandra Interactive Analysis of Observations software package (Fruscione etĀ al., 2006) v4.17 with CALDB version 4.12. Initially, the standard level 1 event lists were reprocessed into the level 2 data products by the chandra_repro script, in which standard grade, status, and good time filters, as well as VFAINT background cleaning were applied. Meanwhile, the routine deflare was utilized to remove background flaring periods. Pileup was estimated using PIMMS version 4.15111https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3pimms/w3pimms.pl. Since B3 1239+376 had % pileup over the 0.5ā7.0 keV energy band, pileup was ignored for this study.
@@
The specextract task was used in the X-ray spectral analysis. The target events were extracted from a circular region with a radius of centered at the position of B3 1239+376, while for background photons nearby circles with a larger radius in the blank region were considered. The numbers of observed net photons of the observation in year 2007 and 2022 are 117 and 198, respectively. We grouped the spectra to have at least five counts per bin. Then the XPEC tool (Arnaud 1996, version 12.14.0h) was used to archive the spectral fitting procedure, in which the the energy range was set as 0.5 ā 7Ā keV and the -statistic (Cash, 1979) was adopted. The spectral model was assumed as a power-law distribution absorbed by the Galactic column density along the line of sight ( = , HI4PI Collaboration etĀ al. 2016). The data together with the optimized models were shown in Figure 4. The analysis on the observation in year 2007 gave an unabsorbed flux of (-Statistic/d.o.f., 12.3/17), along with a hard spectrum . While the results on the other observation indicates a possibly fainter emission, (-Statistic/d.o.f., 26.3/33). The X-ray emission in year 2022 is confirmed to be spectrally hard, .
@@ II.3 SDSS spectrum
During the fit of the SDSS spectrum, effects of galactic extinction was removed firstly, and the continuum was described by a power law and polynomial model plus the Fe ii template (Vestergaard & Wilkes, 2001; Vestergaard & Peterson, 2006). Uncertainties were derived from 1000 mock spectra, generated by adding Gaussian noise consistent with the measurement errors. The standard deviation of the best-fit parameters across these realizations provided the uncertainty for each spectral quantity. As shown in Figure 5, several emission lines are distinct, especially the Ly and the Cāiv lines. The redshift is constrained as =3.82 0.01, consistent with the value list in SDSS quasar catalog DR16 (Wu & Shen, 2022). For the Ly emission line region, we used two components to fit Ly line and one component to fit NV line. The is obtained as , while the is . Taking the bolometric correction factor = 3.81 (Shen etĀ al., 2011), the bolometric disk luminosity is suggested as high as . As for the Cāiv line fitting, three Gaussian components, including one narrow component and two broad components, were adopted. We estimated the as . However, note that the Cāiv line is known to be biased by strong outflows, and hence it should be cautious for virial black hole mass estimations (Dix etĀ al., 2023). Future near-infrared spectroscopy is needed to give a robust BH mass estimation of B3 1239+376 (Burke etĀ al., 2024).
@@ II.4 Infrared data
B3 1239+376 is included in the AllWISE Source Catalog (Wright etĀ al., 2010; WISE Team, 2019), named as WISEA J124209.81+372006.0. The W1 (m) and W2(m) magnitudes listed in the catalog are 16.22 0.06 and 15.90 0.15, respectively. The detections in W3 (m) and W4(m) bands are not statistically significant. Long-term infrared light curves in W1 and W2 bands were extracted, based on the time-resolved WISE/NEOWISE co-added imagesĀ (Meisner etĀ al., 2018), due to the faint infrared radiations. During the analyses, point spread function photometry was carried out, targeting on the the difference images that the first epoch observations were treated as reference images (Jiang etĀ al., 2021). Along with the long-term -ray light curve, the infrared light curves are also shown in Figure 3. A -test with a constant flux level as the null hypothesis was utilized to investigate the significance of the infrared variations. The minimum reduced values obtained by varying the constant flux levels in W1 and W2 light curves are 7.2 and 3.9 (d.o.f. = 23), respectively. Therefore, the infrared variability is suggested to be significant (). The variation tendency in the two bands are similar. The variability amplitude in W2 band appears to be higher than that in the W1 band, but the former also has larger measure uncertainties. At the start of the WISE survey, B3 1239+376 was in the quiescent state, after year 2014, a flux rise was apparent lasting about 1500 days. The detected flux maxima were at MJD 58260, then a flux decline was followed. The emissions bounced back with a peak at MJD 58832. There is a minor flux activity later around MJD 59562.
Although the WISE mission was ended at 31th July 2024 (i.e., MJD 60522), a new infrared space telescope, named as Spectro-Photometer for the History of the Universe, Epoch of Reionization and Ices Explorer (SPHEREx, Bock et al. 2025), has been successfully launched in Mar. 2025. It performs a 0.75-5 µm spectroscopy survey over the entire sky. Based on its first reprocessed quick release products (SPHEREx Team, 2025), one spectrum on B3 1239+376 observed from MJD 60800 to MJD 60829 is obtained. To compare with the WISE measurements, we convolved the SPHEREx spectrum with the WISE W1 and W2 filters transmission curves222https://svo2.cab.inta-csic.es/svo/theory/fps3/index.php. The obtained SPHEREx W1 and W2 flux densities are given as (0.259 0.008) mJy and (0.278 0.32) mJy, respectively, see Figure 3. Interestingly, these values are unprecedentedly high, roughly 2-3 times of majority of the WISE estimations.
@@ II.5 Radio data
B3 1239+376 is a well studied radio source. It was monitored by RATAN-600 (Parijskij, 1993), in 12 epochs between MJD 57836 and 59026, at multi-bands from 1.2Ā GHz to 22Ā GHz (Sotnikova etĀ al., 2021). In addition, seven snapshots at K-band (24Ā GHz) of Very Long Baseline Array (VLBA) observations with submilliarcsecond resolution were carried out (de Witt etĀ al., 2023). Images from the Very Large Array Sky Survey (VLASS, Lacy etĀ al. 2020) were also looked into, from which radio fluxes at 3Ā GHz were extracted by modelling the source with a Gaussian fit on the image plane using the Common Astronomy Software Applications. The VLASS flux density in the first epoch was obtained as 612 9 mJy, which is consistent with the value (i. e., 608 3 mJy) in epoch 1 quick look catalog of VLASS (Gordon etĀ al., 2020). The radio data are also plotted in Figure 3. The RATAN-600 multi-bands light curves exhibit a general behavior of flux decay in long-timescale, which is supported by the results of the VLASS observations. Two components, including a compact core as well as a parsec-scale jet, are resolved in the VLBA images, interestingly, from which different temporal behaviors are revealed. For the core, the K-band flux density is 0.099 Jy at MJD 57807 and then reaches to 0.153 Jy at MJD 58180, with an average relative uncertainty of 9%, indicating a likely flux increase. On the other hand, no significant variation of the jet flux density is found.
@@ II.6 Implications from the multi-wavelength observations
Through a detailed investigation of the multi-wavelength properties of B3 1239+376, its nature as a high- -ray-emitting blazar is revealed. In addition to its typical blazar behavior recorded in literature, for instance, the extremely high radio loudness, super-luminous motion and hard X-ray spectrum (Britzen etĀ al., 2008; McKeough etĀ al., 2016), in this study, -ray signals detected by Fermi-LAT toward B3 1239+376 are reported. In year 2025, a robust (7.7 in global) -ray source there emerges, with a relatively soft spectrum ( = 2.6). The half-year averaged photon flux then is one order of magnitude of that from the initial 9-year Fermi-LAT survey, indicating a strong -ray flare. B3 1239+376 is known as the unique blazar (candidate) in the direction and the association probability is calculated as 0.91, which establishes a physical connection between the -ray source and B3 1239+376. Moreover, by revisiting the Fermi-LAT data, prior -ray activities linking to the target are also found. Although these two detections are marginal (TS 15) each alone, a joint analysis brings enhanced statistics (TS = 31) enabling us to perform a localization analysis, from which B3 1239+376 also falls into the localization uncertainty, shown in Figure 2. Most importantly, behaviors of the -ray and infrared light curves are rather similar. At a scope of 16-year monitoring, -ray flux reaches up to a maximum level in year 2025, at the same time, the infrared fluxes are unprecedentedly high. In addition, the appearance of the prior tentative -ray signals is temporally coincident with the times of the high flux levels in the WISE light curves, see Figure 3. The coincidence acts as the decisive evidence to pin down the association relationship. The most energetic -ray photon relative to B3 1239+376 is at energy of 1Ā GeV, which could not challenge the current EBL models (e.g., Finke etĀ al., 2010). Although the data sampling of the WISE observations is limited, detections of contemporaneous brightening in the Fermi-LAT and infrared light curves play an important role in identification of high- -ray-emitting blazars (Liao etĀ al., 2019; Li etĀ al., 2020). It is also worth mentioning that the k-band core flux climbs when the -ray and infrared emissions brighten. Unfortunately, no continual VLBA monitoring after the activity is available and detecting a potential ejection of a new jet blob could not be accomplished then.
High-redshift (z 3) -ray blazars are rare. Among the thousands of extragalactic -ray emitters (Ajello etĀ al., 2022), only a handful of these sources are found (e.g., Ackermann etĀ al., 2017). Yet they are crucial for our understanding in AGN jet cosmic evolution, the star formation history of the universe, and the initial mass growth of SMBHs. In particular, besides the two sources above of 4, B3 1428+422 and GB 1508+5714 (Ackermann etĀ al., 2017; Liao etĀ al., 2018), there are five known blazars with 3.5 detected by Fermi-LAT to date. They are NVSS J121915+365718 ( = 3.53, Paliya etĀ al. 2020), PMN J2219-2719 ( = 3.63, Li etĀ al. 2020), PKS 0201+113 ( = 3.64, Lei etĀ al. 2024), MG3 J163554+3629 ( = 3.65, Ackermann etĀ al. 2017) and PKS 1351-018 ( = 3.72, Punsly etĀ al. 2021). Therefore, as an identified -ray emitter, B3 1239+376 ( = 3.82) is proposed as the third most distant -ray-emitting blazar so far. Most of these extreme sources are FSRQs with soft -ray spectra, but NVSS J121915+365718 is suggested as a BL Lac (Paliya etĀ al., 2020). B3 1239+376, along with other high- -ray sources, are plotted in a - diagram (i.e., Figure 6). The information listed in 4LAC-DR3 (Ajello etĀ al., 2022), as well as those reported in individual studies, are adopted. Considering the high apparent luminosity (up to in the half-year high flux state) and relatively soft spectrum (), it is not surprising that our target lies with the region occupied by the high- sources. In addition to the significant variations of the flux level, the spectral hardening when brighten is indicated for some sources. However, the investigation of spectral variation is not available for B3 1239+376 due to limited statistics in the first 12-year Fermi-LAT survey. Note that several distant -ray sources are not included in 4LAC-DR3, alike of B3 1239+376. Because the rather faint emissions, they are detectable only in the high flux state, while the background is overwhelming for the majority of the time. Therefore, a detailed temporal analysis is crucial to catch the -ray signal from high- blazars.
III Discussions and Summary
@@
The multiwavelength observations not only bring valuable temporal information of B3 1239+376, but also allow us to draw the broadband SEDs, presented in Figure 7. According to the high flux state, quasi-simultaneous data including the -ray spectrum between MJD 58060 and 58242, ZTF g and r band photometry at MJD 58258 as well as the i band one at MJD 58257, WISE W1 and W2 band measurements at MJD 58260, and the RATAN-600 flux densities at MJD 58235 from 2.3 to 22Ā GHz. On the other hand, the low-flux-state SED consists of the Fermi-LAT upper limit from MJD 59065 to 60626, the Chandra snapshot at MJD 59944, the ZTF g band exposure at MJD 59937, and WISE W1 and W2 estimations at MJDĀ 59928. Additionally, -ray and infrared data correspond to the flare event in year 2025 are also presented, although they are not modeled due to the limited data coverage. Un-simulaneous archival data are collected and plotted as backgrounds, such as the radio data collected from NED, the optical photometric data from SDSS and multi-bands infrared data from Infrared Array Camera (IRAC) aboard Spitzer. No significant flux variations are seen in the optical r and g band, in which the accretion disk emission is likely dominant. The big blue bump is described as a standard Shakura & Sunyaev (1973) disk, extending from 3 to 2000, where is the Schwarzschild radius. The total disk luminosity is produced as , in which represents the accretion rate and the accretion efficiency is set as a typical value, 0.1. The accretion disk emission is characterized as a multi-temperature radial profile. At a certain radius , the local temperature is,
@@
where is the Stefan-Boltzmann constant. Since the g band is likely drop-out due to neutral intergalactic medium absorption, it is not considered for the modeling of the disk. It is remarkable that the infrared radiations significantly vary for different flux states. In the low flux state, the W1 and W2 band data points are in alignment with the accretion disk radiations, while the fluxes increase, especially for the W2 band, in the high flux state when the jet contribution is un-ignorable. The jet activity in the infrared bands is also supported by the un-simulaneous IRAC observations (Teplitz etĀ al., 2010). The change of the -ray spectra is also clear. In the 4-year long low flux state, only an upper limit is yielded for Fermi-LAT. While a significant -ray signal are revealed in the half-year period of 2025. Due to the limited statistics, a butterfly plot is drawn for the -ray spectrum of the high-flux-state SED.
@@
where the are indices of the broken power-law distribution, , and are the minimum, breaking and maximum energies of the electrons, respectively, and is the normalization of the particle number density. The transformations of frequency and luminosity between different frames (i.e., the jet comoving frame and the observational frame) are and . In the leptonic radiation scenario, both synchrotron and IC processes are included. For the former, the synchrotron self-absorption (SSA) process is considered. In the IC processes, the KleinNishina effect is taken. Because of the strong Ly line (i.e., 2.5 Hz at the AGN frame) detected in the SDSS spectrum, it is adopted as the external soft photons in the EC process. Assuming the broad-line photons are 10% of the accretion luminosity in a shell-like shape located at a distance cm (e.g., Ghisellini etĀ al., 2011), the corresponding energy density is set as = 2 erg . The radius of the emitting blob is usually constrained by the observed variability timescale in short term, , however, no such information of B3 1239+376 is available. Thus, a typical value of cm is used.

### v2

@@ II.2.1 Chandra data
B3 1239+376 was observed twice by the Advanced CCD Imaging Spectrometer (ACIS, Garmire et al. 2003), on 2007 Mar. 10 (ACIS-7, ObsID: 7871, PI.: C. Cheung) and 2022 Dec. 31 (ACIS-67, ObsID 26906; PI: E. Meyer). The corresponding exposure times are 5.2 and 20.1 ks, respectively. The observations were performed on the pointing mode, events were recorded in the Faint telemetry format and with the Timed Exposure mode. We carried out the data reductions using the Chandra Interactive Analysis of Observations software package (Fruscione et al., 2006) v4.17 with CALDB version 4.12. Initially, we reprocessed the standard level 1 event lists into the level 2 data products by the chandra_repro script, in which standard grade, status, and good time filters, as well as VFAINT background cleaning were applied. Meanwhile, we removed background flaring periods by the routine deflare. We estimated the possible pileup effect using PIMMS version 4.15111https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3pimms/w3pimms.pl. Since B3 1239+376 had % pileup over the 0.5–7.0 keV energy band, pileup was ignored for this study.
@@
We used the specextract task in the X-ray spectral analysis. We extracted the target events from a circular region with a radius of centered at the position of B3 1239+376, while the background spectrum was extracted from two circular regions with radius in blank sky regions. The numbers of observed net photons of the observation in year 2007 and 2022 are 117 and 198, respectively. We grouped the spectra to have at least five counts per bin. Then we used the XSPEC tool (Arnaud 1996, version 12.14.0h) to perform the spectral fitting procedure, in which the energy range was set as 0.5 – 7 keV and the -statistic (Cash, 1979) was adopted. The spectral model was assumed as a power-law distribution absorbed by the Galactic column density along the line of sight ( = , HI4PI Collaboration et al. 2016). The data together with the optimized models were shown in Figure 4. The analysis on the observation in year 2007 gave an unabsorbed flux of (-Statistic/d.o.f., 12.3/17), along with a hard spectrum . While the results on the other observation indicates a possibly fainter emission, (-Statistic/d.o.f., 26.3/33). The X-ray emission in year 2022 is confirmed to be spectrally hard, .
II.2.2 Swift data
There are six snapshots from Neil Gehrels Swift Observatory (Gehrels et al., 2004) in the direction of B3 1239+376. The observational target is RX 1241.4+3722, a nearby Seyfert I galaxy roughly 8.3′ away from the high- quasar. We analyzed the XRT photon counting mode data as well as the UVOT data with the FTOOLS software version 6.33.2. Initially, we performed the event cleaning for the XRT data by the xrtpipeline task using standard quality cuts. Then we utilized the ximage task to search any X-ray sources in the field. Unfortunately, no significant (i.e., signal-to-noise, S/N 3) excesses over the background towards B3 1239+376 were found. For the UVOT images, we carried out aperture photometries by the uvotsource task, with a 5 arcsec circular aperture along with a background extraction in a larger source-free region. One marginal detection (S/N = 3.4, 21.02 0.02 mag), in B-band at MJD 56220, was revealed.
@@ II.3 SDSS spectrum
During the fit of the SDSS spectrum from SDSS data release 16 (Ahumada et al., 2020), we firstly removed effects of galactic extinction, and described the continuum by a power law and polynomial model plus the Fe ii template (Vestergaard and Wilkes, 2001; Vestergaard and Peterson, 2006). We derived the uncertainties from 1000 mock spectra, generated by adding Gaussian noise consistent with the measurement errors. The standard deviation of the best-fit parameters across these realizations provided the uncertainty for each spectral quantity. As shown in Figure 5, several emission lines are distinct, especially the Ly and the C iv lines. The redshift is constrained as =3.82 0.01, consistent with the value listed in SDSS quasar catalog DR16 (Wu and Shen, 2022). For the Ly emission line region, we used two components to fit the Ly line and one component to fit the NV line. The is obtained as , while the is . Taking the bolometric correction factor = 3.81 (Shen et al., 2011), the bolometric disk luminosity is suggested as high as . As for the C iv line fitting, three Gaussian components, including one narrow component and two broad components, were adopted. We estimated the as . However, note that the C iv line is known to be biased by strong outflows, and hence one should be cautious to base black hole mass estimations on this emission line (Dix et al., 2023). Future near-infrared spectroscopy by a ground-based telescope, like Keck/NIRES (Wilson et al., 2004), would give a robust BH mass estimation of B3 1239+376 (Burke et al., 2024).
@@ II.4 Infrared data
B3 1239+376 is included in the AllWISE Source Catalog (Wright et al., 2010; WISE Team, 2019), named as WISEA J124209.81+372006.0. The W1 (m) and W2(m) magnitudes listed in the catalog are 16.22 0.06 and 15.90 0.15, respectively. The detections in W3 (m) and W4(m) bands are not statistically significant. We extracted long-term infrared light curves in W1 and W2 bands , based on the time-resolved WISE/NEOWISE co-added images (Meisner et al., 2018), due to the faint infrared radiation. During the analyses, we carried out point spread function photometry, targeting the difference images that the first epoch observations were treated as reference images (Jiang et al., 2021). Along with the long-term -ray light curve, the infrared light curves are also shown in Figure 3. We utilized a -test with a constant flux level as the null hypothesis, to investigate the significance of the infrared variations. The minimum reduced values obtained by varying the constant flux levels in W1 and W2 light curves are 7.2 and 3.9 (d.o.f. = 23), respectively. Therefore, the infrared variability is suggested to be significant (). The variation tendency in the two bands are similar. The variability amplitude in W2 band appears to be higher than that in the W1 band, but the former also has larger measure uncertainties. At the start of the WISE survey, B3 1239+376 was in the quiescent state, after year 2014, a flux rise was apparent lasting about 1500 days. The detected flux maxima were at MJD 58260, after which a flux decline followed. The emissions bounced back with a peak at MJD 58832. There is a minor flux activity later around MJD 59562.
Although the WISE mission was ended on 31th July 2024 (i.e., MJD 60522), a new infrared space telescope, named as Spectro-Photometer for the History of the Universe, Epoch of Reionization and Ices Explorer (SPHEREx, Bock et al. 2025), has been successfully launched in Mar. 2025. It performs a 0.75-5 µm spectroscopy survey over the entire sky. Based on its first reprocessed quick release products (SPHEREx Team, 2025), we obtained one spectrum on B3 1239+376 observed from MJD 60800 to MJD 60829. To compare with the WISE measurements, we convolved the SPHEREx spectrum with the WISE W1 and W2 filter transmission curves222https://svo2.cab.inta-csic.es/svo/theory/fps3/index.php. We derived the SPHEREx W1 and W2 flux densities as (0.259 0.008) mJy and (0.278 0.032) mJy, respectively, see Figure 3. Interestingly, these values are unprecedentedly high, roughly 2-3 times higher than the average fluxes observed by WISE.
@@ II.5 Radio data
B3 1239+376 is a well studied radio source. It was monitored by RATAN-600 (Parijskij, 1993), in 12 epochs between MJD 57836 and 59026, at multi-bands from 1.2 GHz to 22 GHz (Sotnikova et al., 2021). In addition, seven snapshots at K-band (24 GHz) of Very Long Baseline Array (VLBA) observations (from year 2017 to 2018) with submilliarcsecond resolution were carried out (de Witt et al., 2023). We looked into images from the Very Large Array Sky Survey (VLASS, Lacy et al. 2020), from which radio fluxes at 3 GHz were extracted by modelling the source with a Gaussian fit on the image plane using the Common Astronomy Software Applications. We derived the VLASS flux density in the first epoch as 612 9 mJy, which is consistent with the value (i. e., 608 3 mJy) in epoch 1 quick look catalog of VLASS (Gordon et al., 2020). The radio data are also plotted in Figure 3. The RATAN-600 multi-bands light curves exhibit a general behavior of flux decay in long-timescale, which is supported by the results of the VLASS observations. Two components, including a compact core as well as a parsec-scale jet, are resolved in the VLBA images, interestingly, from which different temporal behaviors are revealed. For the core, the K-band flux density is 0.099 Jy at MJD 57807 and then reaches to 0.153 Jy at MJD 58180, with an average relative uncertainty of 9%, indicating a likely flux increase. On the other hand, no significant variation of the jet flux density is found.
@@ II.6 Implications from the multi-wavelength observations
Through a detailed investigation of the multi-wavelength properties of B3 1239+376, its nature as a high- -ray-emitting blazar is revealed. In addition to its typical blazar behavior recorded in literature, for instance, the extremely high radio loudness, super-luminous motion and hard X-ray spectrum (Britzen et al., 2008; McKeough et al., 2016), in this study, -ray signals detected by Fermi-LAT toward B3 1239+376 are reported. In year 2025, a robust (7.7 in global) -ray source there emerges, with a relatively soft spectrum ( = 2.6). The half-year averaged photon flux then is one order of magnitude of the value that we derived from the initial 9-year Fermi-LAT observations, indicating a strong -ray flare. B3 1239+376 is known as the unique blazar (candidate) in the direction and the association probability is calculated as 0.91, which establishes a physical connection between the -ray source and B3 1239+376. Moreover, by revisiting the Fermi-LAT data, prior -ray activities linking to the target are also found. Although these two detections are marginal (TS 15) each alone, a joint analysis brings enhanced statistics (TS = 31) enabling us to perform a localization analysis, from which B3 1239+376 also falls into the localization uncertainty, shown in Figure 2. Most importantly, behaviors of the -ray and infrared light curves are rather similar. At a scope of 16-year monitoring, -ray flux reaches up to a maximum level in year 2025, at the same time, the infrared fluxes are unprecedentedly high. In addition, the appearance of the prior tentative -ray signals is temporally coincident with the times of the high flux levels in the WISE light curves, see Figure 3. The coincidence acts as the decisive evidence to pin down the association relationship. The IR and -ray observations suggest that in year 2017/2020 B3 1239+376 underwent mild activities while a strong burst came in the first half of year 2025. The most energetic -ray photon relative to B3 1239+376 is at energy of 1 GeV, which could not challenge the current EBL models (e.g., Finke et al., 2010). Although the data sampling of the WISE observations is limited, detections of contemporaneous brightening in the Fermi-LAT and infrared light curves play an important role in identification of high- -ray-emitting blazars (Liao et al., 2019; Li et al., 2020). It is also worth mentioning that the k-band core flux climbs when the -ray and infrared emissions brighten. Unfortunately, no continual VLBA monitoring after the activity is available and detecting a potential ejection of a new jet blob could not be accomplished then.
III SED modeling
@@
The multiwavelength observations not only bring valuable temporal information of B3 1239+376, but also allow us to compile the broadband SEDs, presented in Figure 6. According to the high flux state, quasi-simultaneous data including the -ray spectrum between MJD 58060 and 58242, ZTF g and r band photometry at MJD 58258 as well as the i band one at MJD 58257, WISE W1 and W2 band measurements at MJD 58260, and the RATAN-600 flux densities at MJD 58235 from 2.3 to 22 GHz. On the other hand, the low-flux-state SED consists of the Fermi-LAT upper limit from MJD 59065 to 60626, the Chandra snapshot at MJD 59944, the ZTF g band exposure at MJD 59937, and WISE W1 and W2 flux densities at MJD 59928. Additionally, -ray and infrared data corresponding to the flare event in year 2025 are also presented, although they are not modeled due to the limited data coverage. Non-simultaneous archival data are collected and plotted as backgrounds, such as the radio data collected from NED, the optical photometric data from SDSS and multi-bands infrared data from Infrared Array Camera (IRAC) aboard Spitzer.
Since no significant flux variations are seen in the optical r and g band, the accretion disk emission is likely dominant. The big blue bump is described as a standard Shakura and Sunyaev (1973) disk, extending from 3 to 2000, where is the Schwarzschild radius. The total disk luminosity is produced as , in which represents the accretion rate and the accretion efficiency is set as a typical value, 0.1. The accretion disk emission is characterized as a multi-temperature radial profile. At a certain radius , the local temperature is,
@@
where is the Stefan-Boltzmann constant. Since the g band is likely drop-out due to neutral intergalactic medium absorption, it is not considered for the modeling of the disk. It is remarkable that the infrared radiations significantly vary for different flux states. In the low flux state, the W1 and W2 band data points are in alignment with the accretion disk radiation, while the fluxes increase, especially for the W2 band, in the high flux state when the jet contribution is un-ignorable. The jet activity in the infrared bands is also supported by the non-simultaneous IRAC observations (Teplitz et al., 2010). The change of the -ray spectra is also clear. In the 4-year long low flux state, only an upper limit is yielded for Fermi-LAT. While a significant -ray signal is revealed in the half-year period of 2025. Due to the limited statistics, a butterfly plot is drawn for the -ray spectrum of the high-flux-state SED.
@@
where the are indices of the broken power-law distribution, , and are the minimum, breaking and maximum energies of the electrons, respectively, and is the normalization of the particle number density. The transformations of frequency and luminosity between different frames (i.e., the jet comoving frame and the observational frame) are and , where is the jet Doppler factor. In the leptonic radiation scenario, both synchrotron and IC processes are included. For the former, the synchrotron self-absorption (SSA) process is considered. In the IC processes, the KleinNishina effect is taken into account. Because of the strong Ly line (i.e., 2.5 Hz at the AGN frame) detected in the SDSS spectrum, it is adopted as the external soft photons in the EC process. Assuming the broad-line photons are 10% of the accretion luminosity in a shell-like shape located at a distance cm (e.g., Ghisellini et al., 2011), the corresponding energy density is set as = 2 erg . The radius of the emitting blob is usually constrained by the observed variability timescale in short term, , however, no such information of B3 1239+376 is available. Thus, a typical value of cm is used.
Due to the limited data coverage of the SEDs, we adopted a ‘fit-by-eye’ strategy to carry out the SED modeling. For the high-flux-state SED, the radio data points were not considered in the jet modeling, because they are likely from an extended region rather than the compact blob, and the radio emissions are seriously suffered by the SSA. Meanwhile, the optical emissions are dominated by the accretion disk. The remaining data points are insufficient to conduct a -test. While for the low-flux-level SED, the situation is similar. We fixed majority of the input parameters for the SEDs of different flux states, such as and , , as well as . We only allowed a change of a minimal amount of parameters (i.e., B, and ). Nevertheless, the simple leptonic model can reasonably reproduce the both SEDs, see Figure 6. The corresponding input parameters are listed in Table 1.
IV Discussions and Summary
High-redshift (z 3) -ray blazars are rare. Among the thousands of extragalactic -ray emitters (Ajello et al., 2022), only a handful of these sources are found (e.g., Ackermann et al., 2017). Yet they are crucial for our understanding in AGN jet cosmic evolution, the star formation history of the universe, and the initial mass growth of SMBHs. In particular, besides the two sources above of 4, B3 1428+422 and GB 1508+5714 (Ackermann et al., 2017; Liao et al., 2018), there are five known blazars with 3.5 detected by Fermi-LAT to date. They are NVSS J121915+365718 ( = 3.53, Paliya et al. 2020), PMN J2219-2719 ( = 3.63, Li et al. 2020), PKS 0201+113 ( = 3.64, Lei et al. 2024), MG3 J163554+3629 ( = 3.65, Ackermann et al. 2017) and PKS 1351-018 ( = 3.72, Punsly et al. 2021). Therefore, as an identified -ray emitter, B3 1239+376 ( = 3.82) is proposed as the third most distant -ray-emitting blazar so far. Most of these extreme sources are FSRQs with soft -ray spectra, but NVSS J121915+365718 is suggested as a BL Lac (Paliya et al., 2020). B3 1239+376, along with other high- -ray sources, are plotted in a - diagram (i.e., Figure 7). The information listed in 4LAC-DR3 (Ajello et al., 2022), as well as those reported in individual studies, are adopted. Considering the high apparent luminosity (up to in the half-year high flux state) and relatively soft spectrum (), it is not surprising that our target lies with the region occupied by the high- sources. In addition to the significant variations of the flux level, the spectral hardening when brighten is indicated for some sources. However, the investigation of spectral variation is not available for B3 1239+376 due to limited statistics in the first 12-year Fermi-LAT survey. Note that several distant -ray sources are not included in 4LAC-DR3, alike of B3 1239+376. Because the rather faint emissions, they are detectable only in the high flux state, while the background is overwhelming for the majority of the time. Therefore, a detailed temporal analysis is crucial to catch the -ray signal from high- blazars.